Financial data for Redstone Company for last year appear below:

Redstone Company
Statements of Financial Position
Beginning Balance Ending Balance
Assets:
Cash $ 303,000 $ 337,392
Accounts receivable 169,000 217,000
Inventory 242,000 238,000
Plant and equipment (net) 481,000 407,000
Investment in Balsam Company 256,000 308,000
Land (undeveloped) 280,000 280,000
Total assets $ 1,731,000 $ 1,787,392
Liabilities and owners’ equity:
Accounts payable $ 171,000 $ 188,000
Long-term debt 810,000 810,000
Owners’ equity 750,000 789,392
Total liabilities and owners’ equity $ 1,731,000 $ 1,787,392
Redstone Company
Income statement
Sales $ 2,220,000
Less operating expenses 1,687,200
Net operating income 532,800
Less interest and taxes:
Interest expense $ 103,600
Tax expense 191,808 295,408
Net income $ 237,392


The company paid dividends of $198,000 last year. The "Investment in Balsam Company" on the statement of financial position represents an investment in the stock of another company


Required:

a. Compute the company's margin, turnover, and return on investment for last year. (Round your intermediate calculations and final answers to 2 decimal places.)

Margin%TurnoverReturn on investment%

b. The Board of Directors of Redstone has set a minimum required return of 40%. What was the company's residual income last year? (Round your intermediate calculations and final answers to 2 decimal places.)

Residual income
Answer #1

Part-a

Net Operating Margin
Net Opearting Income/Net sales
532800/2220000*100= 24%
Total operating asset Turnover
Net Sales/Average opearting assets
=2220000/1197196= 1.85
Return On Investment
Net Operating Income/Average Opearting assets*100
=532800/1197196= 44.5%
Net Profit Margin=Net Income/Sales*100
=237392/2220000= 10.69%
Total Asset Turnover
Net Sales/Average Total assets
=2220000/((1731000+1787392)/2)=1.26
Part-b
Residual Income
Net opearting Income-(Minimum required rate*Average opearting assets)
=532800-(0.4*1197196)=53922
* Average Operating Asset:= (1731000+1787392-280000-280000-256000-308000)/2
Note land(underdeveloped) and investment are not operating assets